OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (–1, 8)?

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (anonymous):

m=(8-5)/(-1-2)

OpenStudy (anonymous):

m=3/-3 m=-1

OpenStudy (anonymous):

Then, plug in that m into y=mx+b you can use either set of coordinates.

OpenStudy (anonymous):

5=-1(2)+b

OpenStudy (anonymous):

5=-2+b 7=b

OpenStudy (anonymous):

final answer is y=-1x+7 y=-x+7

OpenStudy (anonymous):

understand?